Please help me solve: 5[1/2+(3/5x5/6)/5/8]
Answer by rfer(16322)

3/5*5/6=15/30=1/2
----------------
(1/2)/(5/8)=
1/2*8/5=8/10=4/5
------------------
1/2+4/5=
5/10+8/10=13/10
---------------
5*13/10=65/10=6 1/2